# Representation Learning for Distributional Perturbation Extrapolation

### v1

@@ 4.1 Identifiability
  3. (iii)
    [sufficient diversity] the matrix $\widetilde{\mathbf{W}}\mathbf{L}\in\mathbb{R}^{{d_{Z}}\times M}$, where $\mathbf{L}\in\mathbb{R}^{K\times M}$ is the matrix with columns $(\mathbf{l}_{1}-\mathbf{l}_{0})$, …, $(\mathbf{l}_{M}-\mathbf{l}_{0})$, has full row rank, i.e., $\mathrm{rank}(\widetilde{\mathbf{W}}\mathbf{L})={d_{Z}}$, or the same holds for $\mathbf{W}\mathbf{L}$.
Then the latent representation and the effects of the observed perturbation combinations relative to $\mathbf{l}_{0}$ (as captured by $\mathbf{W}\mathbf{L}$) are identifiable up to affine transformation in the following sense:
$\forall\mathbf{z}\in\mathbb{R}^{d_{Z}}:\qquad\widetilde{\mathbf{f}}^{-1}\circ\mathbf{f}(\mathbf{z})$ $=\mathbf{A}\mathbf{z}+\mathbf{b},$ (4.2)
$\widetilde{\mathbf{W}}\mathbf{L}$ $=\mathbf{A}\mathbf{W}\mathbf{L},$ (4.3)
where $\mathbf{A}:=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}}$ and $\mathbf{b}:=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}+(\widetilde{\mathbf{W}}-\mathbf{A}\mathbf{W})\mathbf{l}_{0}$.
Corollary 4.2 (Affine recovery of the perturbation matrix).
If, in addition to the assumptions of Thm. 4.1, $\mathbf{L}\in\mathbb{R}^{K\times M}$ has full row rank (i.e., $\mathrm{rank}(\mathbf{L})=K\leq M$), then the perturbation matrix $\mathbf{W}$ is identifiable up to affine transformation in the sense that
$\widetilde{\mathbf{W}}=\mathbf{A}\mathbf{W},$ (4.4)
for $\mathbf{A}:=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}}$. In this case, the expression for $\mathbf{b}$ in (4.2) simplifies to $\mathbf{b}=\widetilde{\bm{\mu}}-\mathbf{A}\bm{\mu}$.
The proofs of Thms. 4.1 and 4.2 are provided in §§ A.1 and A.2.
Discussion.
Thm. 4.1 can be interpreted as follows. Fix a set of perturbation labels $(\mathbf{l}_{0},...,\mathbf{l}_{M})$ and a data generating process parametrised by $(\mathbf{f},\mathbf{W},\bm{\mu},\bm{\Sigma})$. Then, for all $\widetilde{\bm{\mu}}$ and $\widetilde{\bm{\Sigma}}$, and for all $\widetilde{\mathbf{W}}$ such that (4.3) holds for $\mathbf{A}=\widetilde{\bm{\Sigma}}^{\frac{1}{2}}\bm{\Sigma}^{-\frac{1}{2}}$, there exists a unique $\widetilde{\mathbf{f}}$, characterized by (4.2), such that $(\widetilde{\mathbf{f}},\widetilde{\mathbf{W}},\widetilde{\bm{\mu}},\widetilde{\bm{\Sigma}})$ gives rise to the same observed distributions of $\mathbf{X}_{e}$. At the same time, any $(\widetilde{\mathbf{f}},\widetilde{\mathbf{W}},\widetilde{\bm{\mu}},\widetilde{\bm{\Sigma}})$ for which this holds for all $e\in[M]_{0}$ is of that form. In other words, the mean $\bm{\mu}$ and covariance $\bm{\Sigma}$ of the basal state are unidentifiable, but the mixing function $\mathbf{f}$ and the relative shift matrix $\mathbf{W}\mathbf{L}$ are identifiable up to affine transformation—provided that the observed training perturbation conditions are sufficiently diverse, as formalised by assumption (iii).
@@ Discussion.
The matrix product $\mathbf{W}\mathbf{L}\in\mathbb{R}^{{d_{Z}}\times M}$ captures the relative effects of the observed perturbations. Specifically, $(\mathbf{W}\mathbf{L})_{je}$ corresponds to the shift in the $j$th latent $Z_{j}$ resulting from $\mathbf{l}_{e}$, relative to a reference condition $\mathbf{l}_{0}$. Moreover, assumption (iii) of Thm. 4.1 implies
$\min\left\{\mathrm{rank}\left(\widetilde{\mathbf{W}}\right),\mathrm{rank}(\mathbf{L})\right\}\geq\mathrm{rank}(\widetilde{\mathbf{W}}\mathbf{L})={d_{Z}}.$ (4.5)
Hence, sufficient diversity requires at least ${d_{Z}}$ elementary perturbations whose associated shift vectors $\mathbf{w}_{k}\in\mathbb{R}^{d_{Z}}$ are linearly independent, and we must observe at least ${d_{Z}}$ perturbation conditions $\mathbf{l}_{e}$ other than $\mathbf{l}_{0}$ such that the relative perturbation vectors $(\mathbf{l}_{e}-\mathbf{l}_{0})\in\mathbb{R}^{K}$ are linearly independent.
@@
The choice of reference environment is arbitrary. Here, we choose $e=0$ as reference without loss of generality. Intuitively, if a perturbation is always present (e.g., $l_{e,1}=1$ for all $e$), then its effects cannot be discerned from the basal state. Therefore, only the effects of the relative perturbations $(\mathbf{l}_{e}-\mathbf{l}_{0})$ can be recovered. In practice, we often have access to an unperturbed, purely observational control condition with $\mathbf{l}_{0}=\bm{0}$.
@@
The mixing function $\mathbf{f}$ in Thm. 4.1 is deterministic, i.e., does not take a separate noise variable ${\bm{\varepsilon}}$ as input, cf. (3.4). In principle, noise can be appended to $\mathbf{z}^{\mathrm{pert}}$ as additional dimensions that are not influenced by perturbations. However, this increases ${d_{Z}}$ and thus makes it harder to satisfy sufficient diversity (see Remark 4.3). Alternatively, the setting of additive noise, $\mathbf{f}(\mathbf{Z})+{\bm{\varepsilon}}$, can be reduced to the noiseless case (Khemakhem et al., 2020).
@@ 4.2 From Identifiability to Extrapolation
Since we aim to make distributional predictions for new perturbations $\mathbf{l}_{\mathrm{test}}$, identifiability is only of intermediary interest. The following result highlights the usefulness of the affine identifiability established in Thm. 4.1 for extrapolation. In particular, the result allows us to uniquely predict the observable effects of certain unseen perturbations—specifically, those which can be expressed as linear combinations of the observed perturbations.
@@
Under the same setting and assumptions as in Thm. 4.1, let $\mathbf{l}_{\mathrm{test}}\in\mathbb{R}^{K}$ be an unseen perturbation label such that
$\left(\mathbf{l}_{\mathrm{test}}-\mathbf{l}_{0}\right)\in\mathrm{span}\left(\left\{\mathbf{l}_{e}-\mathbf{l}_{0}\right\}_{e\in[M]}\right).$ (4.6)
Then the effect of $\mathbf{l}_{\mathrm{test}}$ is uniquely identifiable in the sense that
[…]
The proof of Thm. 4.6 is provided in § A.3.
@@
For our affine identifiability result (Thm. 4.1), it is not necessary that the mean shifts are of the form $\mathbf{W}\mathbf{l}_{e}$. If we replace $\mathbf{W}\mathbf{l}_{e}$ and $\widetilde{\mathbf{W}}\mathbf{l}_{e}$ in (4.1) with arbitrary shift vectors $\mathbf{c}_{e},\widetilde{\mathbf{c}}_{e}\in\mathbb{R}^{{d_{Z}}}$, the same result can be shown to hold with $\mathbf{W}\mathbf{L}$ and $\widetilde{\mathbf{W}}\mathbf{L}$ replaced by $\mathbf{C}$ and $\widetilde{\mathbf{C}}$, defined as the matrices with columns $(\mathbf{c}_{e}-\mathbf{c}_{0})$ and $(\widetilde{\mathbf{c}}_{e}-\widetilde{\mathbf{c}}_{0})$, respectively. That is, the relative shift vectors are identifiable up to affine transformation, regardless of whether they are linear in $\mathbf{l}$. This has implications, e.g., for the CPA model of Lotfollahi et al. (2023) which includes element-wise nonlinear dose-response functions applied to $\mathbf{l}$, see § B.4 for details. However, linearity is leveraged in the proof of our extrapolation result (Thm. 4.6) where it is used in (4.6) to establish a link between $\mathbf{l}_{\mathrm{test}}$ and the training perturbations. Since only $\mathbf{l}_{\mathrm{test}}$ is observed at test time, the above argument thus cannot easily be extended to the extrapolation setting, as this would require establishing a link between $\mathbf{c}_{\mathrm{test}}$ and the training shifts, all of which are unobserved.
Figure 2: Overview of the Perturbation Distribution Autoencoder (PDAE). In a PDAE, the distribution of a target perturbation condition $h$ (purple) is simulated by encoding, perturbing, and (stochastically) decoding data from a source condition $e$ (blue). Dashed arrows indicate model inputs and green boxes model components with trainable parameters. The learning objective (orange, dotted arrows) consists of: a perturbation loss (bottom), measuring, for all pairs of training domains $(e,h)$, the dissimilarity between the true distribution of $\mathbf{X}_{h}$ and its simulated version based on domain $e$; and a distributional reconstruction loss (top), measuring, for each source domain $e$, the dissimilarity between the true conditional distribution of source observations $\mathbf{X}_{e}$ that are mapped to the same encoding $\widehat{\mathbf{Z}}_{e}$ and the corresponding distribution induced by the stochastic decoder. At test time, the target perturbation label $\mathbf{l}_{h}$ can be replaced with an unseen  $\mathbf{l}_{\mathrm{test}}$ to make predictions.
5 Estimation and Prediction: The Perturbation Distribution Autoencoder (PDAE)
@@
To leverage the extrapolation guarantees of Thm. 4.6, we seek to estimate the identifiable parts of the generative process by matching the distributions of the observed data in (2.1) across all perturbation conditions, as prescribed by condition (4.1) in Thm. 4.1. To this end—rather than fitting a probabilistic model by maximising (a lower bound to) the likelihood, as in most existing methods—we take inspiration from the distributional principal autoencoder (Shen & Meinshausen, 2024a) and adapt it for multi-domain perturbation modelling. Our method, the perturbation distribution autoencoder (PDAE), comprises an encoder, a perturbation matrix, and a (stochastic) decoder, which are trained to maximise the similarity between pairs of true and simulated perturbation distributions, see Fig. 2 for an overview. Intuitively, instead of specifying a (Gaussian, as per Thm. 4.1) prior on the basal state from which to sample latents as in (3.1), we use encoded observations from a source domain as perturbed latents and transform them to perturbed latents from a target domain using the learnt perturbation matrix and the assumed additivity as in (3.2).
5.1 Model Components and Outputs
Encoder.
@@ 5.1 Model Components and Outputs
The encoder $\widehat{\mathbf{g}}:\mathbb{R}^{{d_{X}}}\to\mathbb{R}^{{d_{Z}}}$ maps observations $\mathbf{x}$ to the space of perturbation-relevant latents $\mathbf{z}$. Ideally, it should invert the stochastic mixing function in (3.4) in the sense of recovering the perturbed latent state $\mathbf{z}^{\mathrm{pert}}_{e,i}$ (realization of $\mathbf{Z}^{\mathrm{pert}}_{e,i}$ in (3.2)) from observation $\mathbf{x}_{e,i}$ (realization of $\mathbf{X}_{e,i}$). [^2] This is one of the key differences to other methods like CPA (Lotfollahi et al., 2023) and SAMS-VAE (Bereket & Karaletsos, 2023), which seek an encoder that maps to the latent basal state, regardless of the domain $e$. We therefore denote the encoder outputs by
@@
If the encoder recovers the perturbed latents up to affine transformation, the additivity of perturbation effects assumed in (3.2) allows us to map between the latent distributions underlying different perturbation conditions. Specifically, we use a perturbation model parametrised by a perturbation matrix $\widehat{\mathbf{W}}\in\mathbb{R}^{{d_{Z}}\times K}$ to create synthetic perturbed latents from domain $h$ as:
$\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to h,i}:=\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}-\widehat{\mathbf{W}}\mathbf{l}_{e}+\widehat{\mathbf{W}}\mathbf{l}_{h}=\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}+\widehat{\mathbf{W}}(\mathbf{l}_{h}-\mathbf{l}_{e}).$ (5.2)
@@
which can be interpreted as undoing the effects of perturbation $\mathbf{l}_{e}$ (i.e., mapping back to the latent basal state) and then simulating perturbation $\mathbf{l}_{h}$. (For $h=e$, this has no effect and $\mathbf{z}^{\mathrm{pert}}_{e\to h,i}=\mathbf{z}^{\mathrm{pert}}_{e,i}$.)
Decoder.
@@
The decoder $\widehat{\mathbf{f}}:\mathbb{R}^{{d_{Z}}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{{d_{X}}}$ maps estimated perturbed latents $\widehat{\mathbf{z}}^{\mathrm{pert}}$ and noise ${\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$ to observations. When evaluated on synthetic perturbed latents from (5.2), we refer to the corresponding (random) outputs
@@
$\widehat{\mathbf{X}}_{e\to h,i}=\widehat{\mathbf{f}}\left(\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to h,i},\,{\bm{\varepsilon}}\right)\qquad\text{where}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$ (5.3)
as synthetic observations (from domain $h$ based on domain $e$). For $h=e$, i.e., when evaluated on the estimated perturbed latents from (5.1), we refer to the corresponding (random) outputs as reconstructions,
$\widehat{\mathbf{X}}_{e,i}=\widehat{\mathbf{f}}\left(\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i},\,{\bm{\varepsilon}}\right)\qquad\text{where}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}.$ (5.4)
5.2 Induced Distributions
Conditional Distribution Over Reconstructions.
For any given observation $\mathbf{x}$, we can sample reconstructions from the stochastic decoder by fixing its first argument to the encoding $\widehat{\mathbf{g}}(\mathbf{x})$ and sampling noise from $\mathbb{Q}_{\bm{\varepsilon}}$ as in (5.4). Together, the encoder, decoder, and noise distribution thus induce a distribution over reconstructions, formally given by the push-forward $\widehat{\mathbf{f}}\left(\widehat{\mathbf{g}}(\mathbf{x}),\cdot\right)_{\#}\mathbb{Q}_{\bm{\varepsilon}}$. Since this distribution should ideally match that of $\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{x})$ (i.e., the true distribution of observations that share the same encoding), [^3] For example, if $\widehat{\mathbf{g}}$ is injective, the true reconstruction distribution is a point mass on $\mathbf{x}$, whereas if $\widehat{\mathbf{g}}$ is constant, it is the full (unconditional) distribution of $\mathbf{X}$. In this sense, the distribution of reconstructions captures the information about $\mathbf{x}$ that is left unexplained given only its encoding (Shen & Meinshausen, 2024a). we denote it by ^$\widehat{\mathbb{P}}_{\mathbf{X}|\widehat{\mathbf{g}}(\mathbf{X})=\widehat{\mathbf{g}}(\mathbf{x})}$. (Here, we use the notation $\widehat{\mathbb{P}}$ to denote a distribution induced by our model; it does not refer to an empirical distribution.)
Marginal Perturbation Distributions.
@@ 5.2 Induced Distributions
Given a distribution $\mathbb{P}_{e}$ and the corresponding perturbation label $\mathbf{l}_{e}$, our model facilitates sampling synthetic observations for another perturbation condition with label $\mathbf{l}_{h}$ via (5.2) and (5.3). We denote the resulting distribution by ^$\widehat{\mathbb{P}}_{e\to h}$, formally defined as the distribution of
$\widehat{\mathbf{f}}\left(\widehat{\mathbf{g}}\left(\mathbf{X}_{e}\right)+\widehat{\mathbf{W}}\left(\mathbf{l}_{h}-\mathbf{l}_{e}\right),{\bm{\varepsilon}}\right)\qquad\text{where}\qquad\mathbf{X}_{e}\sim\mathbb{P}_{e}\qquad\text{and}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}.$ (5.5)
@@
Crucially, both $\mathbf{X}_{e}$ and ${\bm{\varepsilon}}$ are random in (5.5). This contrasts with the conditional distribution over reconstructions, for which the first argument of $\widehat{\mathbf{f}}$ is considered fixed and the only stochasticity stems from the noise ${\bm{\varepsilon}}$. Intuitively, when simulating observations from $\mathbb{P}_{h}$ based on $\mathbb{P}_{e}$ via (5.5), the random variable $\widehat{\mathbf{g}}(\mathbf{X}_{e})-\widehat{\mathbf{W}}\mathbf{l}_{e}$ in the first argument of $\widehat{\mathbf{f}}$ plays the role of a random latent basal state from (3.1).
5.3 Population-Level Learning Objective
@@
At the population level (i.e., assuming access to the true distributions $\mathbb{P}_{e}$), we propose the following learning objective:
$\left(\widehat{\mathbf{g}}^{*},\widehat{\mathbf{f}}^{*},\widehat{\mathbf{W}}^{*}\right)\in\operatorname*{arg\,min}_{\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}}}\,\mathcal{L}_{\mathrm{pert}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{l}_{e}\right)\right\}_{e\in[M]_{0}}\right)+\lambda\mathcal{L}_{\text{cond-rec}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}};\left\{\mathbb{P}_{e}\right\}_{e\in[M]_{0}}\right),$ (5.6)
where $\mathcal{L}_{\mathrm{pert}}$ is a perturbation loss, which depends on the full model and perturbation labels; $\mathcal{L}_{\text{cond-rec}}$ is a (conditional) reconstruction loss, which only depends on the encoder, decoder, and observed distributions and which may be viewed as a regulariser that is helpful in practice; and $\lambda\geq 0$ is a hyperparameter, which trades off the two losses.
Perturbation Loss.
The perturbation loss $\mathcal{L}_{\mathrm{pert}}$ takes the form of the following sum of pairwise distribution losses
$\mathcal{L}_{\mathrm{pert}}\left(\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}};\left\{\left(\mathbb{P}_{e},\mathbf{l}_{e}\right)\right\}_{e\in[M]_{0}}\right)$ $:=\sum_{e,h\in[M]_{0}}d\left(\widehat{\mathbb{P}}_{e\to h},\mathbb{P}_{h}\right)$ (5.7)

### v2

@@ 4.1 Identifiability
Remark 4.3 (Choice of base distribution).
The choice of means and covariances in assumption (ii) of Thm.˜4.1 serves to eliminate inherent ambiguities of the model class due to overparametrisation. As shown by Lemma˜A.1, this comes w.l.o.g., in the sense that any model with $\mathbb{P}=\mathcal{N}(\bm{\mu},\bm{\Sigma})$ can be reparametrised to take the form in assumption (ii).
@@
The matrix product $\mathbf{W}\mathbf{A}\in\mathbb{R}^{{d_{Z}}\times M}$ captures the relative effects of the observed perturbations since $(\mathbf{W}\mathbf{A})_{je}$ corresponds to the shift in the $j$th latent $Z_{j}$ resulting from $\mathbf{a}_{e}$, relative to a reference condition $\mathbf{a}_{0}$. Moreover, assumption (iii) of Thm.˜4.1 implies
$\min\big\{\mathrm{rank}\big(\widetilde{\mathbf{W}}\big),\mathrm{rank}(\mathbf{A})\big\}\geq\mathrm{rank}(\widetilde{\mathbf{W}}\mathbf{A})={d_{Z}}.$ (4.4)
Hence, sufficient diversity requires at least $d_{Z}$} elementary perturbations whose associated shift vectors $\mathbf{w}_{k}\in\mathbb{R}^{d_{Z}}$ are linearly independent, and we must observe at least $d_{Z}$} perturbation conditions $\mathbf{a}_{e}$ other than $\mathbf{a}_{0}$ such that the relative perturbation vectors $(\mathbf{a}_{e}-\mathbf{a}_{0})\in\mathbb{R}^{K}$ are linearly independent.
@@
The choice of reference environment is arbitrary. Here, we choose $e=0$ as reference without loss of generality. Intuitively, if a perturbation is always present (e.g., $a_{e,1}=1$ for all $e$), then its effects cannot be discerned from the basal state. Therefore, only the effects of the relative perturbations $(\mathbf{a}_{e}-\mathbf{a}_{0})$ can be recovered. In practice, we often have access to an unperturbed, purely observational control condition with $\mathbf{a}_{0}=\bm{0}$.
@@
The mixing function $\mathbf{f}$ in Thm.˜4.1 is assumed deterministic, i.e., does not take noise $\bm{\varepsilon}$} as input, cf. (3.2). In principle, this does not pose a restriction since noise can be appended to $\mathbf{z}^{\mathrm{pert}}$ as additional dimensions that are not influenced by perturbations. However, this increases $d_{Z}$} and makes it harder to satisfy sufficient diversity (see Remark˜4.4). Alternatively, the additive noise setting, $\mathbf{X}=\mathbf{f}(\mathbf{Z})+{\bm{\varepsilon}}$, can be reduced to the noiseless case (Khemakhem et al., 2020).
Figure 2: Overview of the perturbation distribution autoencoder (PDAE). A PDAE simulates the distribution of a target perturbation condition $h$ (purple) by encoding, perturbing, and (stochastically) decoding data from a source condition $e$ (blue). Dashed arrows indicate model inputs and green boxes model components with trainable parameters. The learning objective (orange, dotted arrows) amounts to a perturbation loss (bottom), measuring, for all pairs of training domains $(e,h)$, the dissimilarity between the true distribution of $\mathbf{X}_{h}$ and its simulated version based on domain $e$; At test time, the target perturbation label $\mathbf{a}_{h}$ can be replaced with an unseen $\mathbf{a}_{\mathrm{test}}$ to make predictions.
@@ 4.2 From identifiability to extrapolation
Since we aim to make distributional predictions for new perturbations $\mathbf{a}_{\mathrm{test}}$, identifiability is only of intermediary interest. The following result highlights the usefulness of identifiability up to orthogonal transformation established in Thm.˜4.1 for extrapolation.
It allows us to uniquely predict the observable effects of certain unseen perturbations—those which can be expressed as linear combinations of the observed perturbations, as illustrated in the inset figure.
@@
Under the same setting and assumptions as in Thm.˜4.1, let $\mathbf{a}_{\mathrm{test}}\in\mathbb{R}^{K}$ be an unseen perturbation s.t.
$\left(\mathbf{a}_{\mathrm{test}}-\mathbf{a}_{0}\right)\in\mathrm{span}\big(\left\{\mathbf{a}_{e}-\mathbf{a}_{0}\right\}_{e\in[M]}\big).$ (4.5)
Then the effect of $\mathbf{a}_{\mathrm{test}}$ is uniquely identifiable in that
$\mathbf{X}_{\mathrm{test}}=\mathbf{f}(\mathbf{Z}+\mathbf{W}\mathbf{a}_{\mathrm{test}})\overset{d}{=}\widetilde{\mathbf{f}}\big(\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{a}_{\mathrm{test}}\big)$ (4.6)
@@
For our identifiability result it is not necessary that the mean shifts are of the form $\mathbf{W}\mathbf{a}_{e}$. If we replace $\mathbf{W}\mathbf{a}_{e}$ and $\widetilde{\mathbf{W}}\mathbf{a}_{e}$ in (4.1) with arbitrary shift vectors $\mathbf{c}_{e},\widetilde{\mathbf{c}}_{e}\in\mathbb{R}^{{d_{Z}}}$, the same result as Thm.˜4.1 can be shown to hold with $\mathbf{W}\mathbf{A}$ and $\widetilde{\mathbf{W}}\mathbf{A}$ replaced by $\mathbf{C}$ and $\widetilde{\mathbf{C}}$, defined as the matrices with columns $(\mathbf{c}_{e}-\mathbf{c}_{0})$ and $(\widetilde{\mathbf{c}}_{e}-\widetilde{\mathbf{c}}_{0})$, respectively. That is, the relative shift vectors are identifiable (up to orthogonal transformation), regardless of whether they are linear in $\mathbf{a}$. This is relevant, for example, when including nonlinear interactions or learnable element-wise nonlinear dose-response functions $h_{j}(a_{j})$ as in CPA (Lotfollahi et al., 2023). However, linearity is leveraged in the proof of our extrapolation result, where it is used in (4.5) to establish a link between $\mathbf{a}_{\mathrm{test}}$ and the training perturbations. Since only $\mathbf{a}_{\mathrm{test}}$ is observed at test time, Thm.˜4.7 cannot easily be extended to arbitrary shifts, as this would require establishing a link between $\mathbf{c}_{\mathrm{test}}$ and the training shifts, all of which are unobserved.
4.3 Connection to causal representation learning
In causal inference, experimental data resulting from perturbations is modeled via interventions in an underlying causal model. In structural equation models (SEMs; Pearl, 2009; Bongers et al., 2021), interventions modify a subset of assignments that determine each variable from its direct causes and unexplained noise. In general, our model for the effect of perturbations in latent space (§˜3) differs from how interventions are treated in SEMs. However, as summarized in the following result, the class of shift interventions in linear SEMs constitutes a special case of our model class.
Proposition 4.9 (Modeling shift interventions in SEMs).
Consider a linear SEM, $\mathbf{Z}:=\mathbf{B}^{\top}\mathbf{Z}+\bm{\eta}$, with noise $\bm{\eta}\sim\mathbb{P}_{\bm{\eta}}$ and $\mathbf{B}\in\mathbb{R}^{{d_{Z}}\times{d_{Z}}}$ a weighted adjacency matrix with spectral radius (largest absolute eigenvalue) $\rho(\mathbf{B})<1$. Then, the distribution induced by shift interventions $\mathbf{a}_{e}\in\mathbb{R}^{{d_{Z}}}$,
$\mathbf{Z}:=\mathbf{B}^{\top}\mathbf{Z}+\bm{\eta}+\mathbf{a}_{e},$ (4.7)
is identical to that induced by (3.1) with $\mathbf{W}=(\mathbf{I}-\mathbf{B}^{\top})^{-1}$ and basal state distribution $\mathbb{P}_{\mathbf{Z}}=(\mathbf{I}-\mathbf{B}^{\top})^{-1}_{\#}\mathbb{P}_{\bm{\eta}}$.
Remark 4.10 (Applicability to cyclic or confounded SEMs).
Prop.˜4.9 does not require the acyclicity or unconfoundedness (i.e., independence of the noise components) of the SEM. The spectral radius condition (which holds for acyclic systems but is strictly weaker) ensures stability and invertibility of $\mathbf{I}-\mathbf{B}^{\top}$ (Zheng et al., 2018; Brown et al., 2025).
Combined with a decoder from (3.2), Prop.˜4.9 highlights in which sense our model can be interpreted as interventional causal representation learning (Schölkopf et al., 2021; von Kügelgen et al., 2023; Varici et al., 2025) with a linear latent causal model (Squires et al., 2023; Buchholz et al., 2024) and generalized shift interventions (Zhang et al., 2024). For $K={d_{Z}}$ elementary perturbations, the adjacency matrix corresponding to an inferred $\widehat{\mathbf{W}}$ is given by $\widehat{\mathbf{B}}=(\mathbf{I}-\widehat{\mathbf{W}}^{-1})^{\top}$. However, due to the rotation ambiguity, additional assumptions such as sparse (e.g., single node) shifts would be needed to recover the causal graph induced by the true $\mathbf{B}$.
5 Perturbation distribution autoencoder
@@
To leverage the extrapolation guarantees of Thm.˜4.7, we seek to estimate the identifiable parts of the generative process by matching the distributions of the observed data in (2.1) across all perturbation conditions, as prescribed by condition (4.1) in Thm.˜4.1. To this end, we take inspiration from prior distributional learning approaches (Bouchacourt et al., 2016; Shen & Meinshausen, 2024a) and adapt them for multi-domain perturbation modeling. Our method, the perturbation distribution autoencoder (PDAE), comprises an encoder, a perturbation matrix, and a (stochastic) decoder, which are trained to maximise the similarity between pairs of true and simulated perturbation distributions, see Fig.˜2 for an overview. Intuitively, instead of specifying a (Gaussian, as per Thm.˜4.1) prior on the basal state from which to sample latents as in (3.1), we use encoded observations from a source domain as perturbed latents and transform them to perturbed latents from a target domain using the learnt perturbation matrix and the assumed additivity as in (3.1).
@@
Encoder. The encoder $\widehat{\mathbf{g}}:\mathbb{R}^{{d_{X}}}\to\mathbb{R}^{{d_{Z}}}$ maps observations $\mathbf{x}$ to the space of perturbation-relevant latents $\mathbf{z}$. Ideally, it should invert the stochastic mixing function in (3.2) in the sense of recovering the perturbed latent state $\mathbf{z}^{\mathrm{pert}}_{e,i}$from observation $\mathbf{x}_{e,i}$. [^1] This is a key difference to CPA (Lotfollahi et al., 2023) and SAMS-VAE (Bereket & Karaletsos, 2023), which seek an encoder that maps to the latent basal state, regardless of the domain $e$. We therefore denote the encoder outputs by
@@
Perturbation module. If the encoder recovers the perturbed latents up to orthogonal transformation, the additivity of perturbation effects assumed in (3.1) allows us to map between the latent distributions underlying different perturbation conditions. Specifically, we use a perturbation model parametrised by a perturbation matrix $\widehat{\mathbf{W}}\in\mathbb{R}^{{d_{Z}}\times K}$ to create synthetic perturbed latents from domain $h$ as:
$\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to h,i}:=\widehat{\mathbf{z}}^{\mathrm{pert}}_{e,i}+\widehat{\mathbf{W}}(\mathbf{a}_{h}-\mathbf{a}_{e}).$ (5.2)
@@
which can be interpreted as undoing the effect of $\mathbf{a}_{e}$ (i.e., mapping back to the basal state) and then simulating $\mathbf{a}_{h}$.
@@
Decoder. The decoder $\widehat{\mathbf{f}}:\mathbb{R}^{{d_{Z}}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{{d_{X}}}$ maps estimated perturbed latents $\widehat{\mathbf{z}}^{\mathrm{pert}}$ and noise ${\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$ to observations. When evaluated on synthetic perturbed latents from (5.2), we refer to the corresponding (random) outputs
@@
$\widehat{\mathbf{X}}_{e\to h,i}=\widehat{\mathbf{f}}\big(\widehat{\mathbf{z}}^{\mathrm{pert}}_{e\to h,i},\,{\bm{\varepsilon}}\big)\qquad\text{where}\qquad{\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$ (5.3)
as synthetic observations from domain $h$ based on domain $e$.
@@
Simulated perturbation distributions. Given a distribution $\mathbb{P}_{e}$ and the corresponding perturbation label $\mathbf{a}_{e}$, our model facilitates sampling synthetic observations for another perturbation condition with label $\mathbf{a}_{h}$ via (5.2) and (5.3). We denote the corresponding distribution by $\widehat{\mathbb{P}}_{e\to h}$, formally defined as the distribution of
$\widehat{\mathbf{f}}\big(\widehat{\mathbf{g}}(\mathbf{X}_{e})+\widehat{\mathbf{W}}(\mathbf{a}_{h}-\mathbf{a}_{e}),{\bm{\varepsilon}}\big),$ (5.4)
@@
where both $\mathbf{X}_{e}\sim\mathbb{P}_{e}$ and ${\bm{\varepsilon}}\sim\mathbb{Q}_{\bm{\varepsilon}}$ are random. Intuitively, when simulating $\mathbb{P}_{h}$ based on $\mathbb{P}_{e}$ via (5.4), the random variable $\widehat{\mathbf{g}}(\mathbf{X}_{e})-\widehat{\mathbf{W}}\mathbf{a}_{e}$} in the first argument of $\widehat{\mathbf{f}}$ plays the role of a random latent basal state from (3.1).
@@
Population-level learning objective. At the population level (i.e., assuming access to the true distributions $\mathbb{P}_{e}$), we propose minimizing a sum of pairwise distribution losses,
$\left(\widehat{\mathbf{g}}^{*},\widehat{\mathbf{f}}^{*},\widehat{\mathbf{W}}^{*}\right)\in\operatorname*{arg\,min}_{\widehat{\mathbf{g}},\widehat{\mathbf{f}},\widehat{\mathbf{W}}}\,\sum_{e,h\in[M]_{0}}d\left(\widehat{\mathbb{P}}_{e\to h},\mathbb{P}_{h}\right)\vskip-2.5pt$ (5.5)